The temperature coefficient of resistance for carbon at 0oC is -0.0048/oC. What is the significance of the minus sign? A carbon resistor has a resistance of 500Ω.Determine its resistance at 50oC
1
Expert's answer
2020-05-07T10:01:12-0400

Minus sign shows that on increasing the temperature the resistance value will decrease or vice versa


"R=R_o(1+\alpha(T-T_o)"

"R=500(1+(-0.0048(50-0))=500(1-0.0048\times50)=500\times0.76=380 ohm"